OpenStudy (anonymous):

multiply and write the answer without exponents (2*10 4)*(4*10 5)

OpenStudy (anonymous):

it's all multiplication so you can rearrange using the commutative property 4*4*2*5*10*10 = 4*4*10*10*10 = 16*10*10*10 and for each 10, add a 0 to the end of 16

OpenStudy (anonymous):

So it wouldnt be 0.8??

OpenStudy (anonymous):

? is there a division somewhere?

OpenStudy (anonymous):

No its not im just confused about the answer if its 0.8 or 8, 000, 000, 000

OpenStudy (anonymous):

please rewrite using equation editor because I'm not sure what your question actually is

OpenStudy (anonymous):

(2*10 to the 4th power)*(4*10 to the 5th power)

OpenStudy (anonymous):

add the exponents to get 9. there's that many 0 after 2*4 which is 8

OpenStudy (anonymous):

Thanks